# 2M-Belebele: Highly Multilingual Speech and American Sign Language Comprehension Dataset

### v1

@@ Abstract
We introduce the first highly multilingual speech and American Sign Language (ASL) comprehension dataset by extending Belebele. Our dataset covers 74 spoken languages at the intersection of Belebele and FLEURS, and one sign language (ASL). We evaluate 2M-Belebele dataset for both 5-shot and zero-shot settings and across languages, the speech comprehension accuracy is 8% average lower compared to reading comprehension.
@@ 1 Introduction
The recent Belebele benchmark is the first corpus that addresses text reading comprehension for a large number of languages following a multi-way parallel approach (Bandarkar et al., 2023). The entire Belebele text statistics are summarized in Table 1. Currently, there are no highly multilingual evaluation datasets for natural language understanding (NLU) that cover either both speech and text, and/or ASL.
The outstanding performance of some MT and text-to-speech (TTS) models has enabled a rise in the number of works using synthetically generated training data. Furthermore, some recent works propose to also use synthetic data for evaluation; e.g., Üstün et al. (2024); Seamless-Communication et al. (2023); Nguyen et al. (2024); Nachmani et al. (2023). This strategy allows researchers to extend datasets to low-resource languages and to other modalities, such as speech. However, we prove that using synthetic data for evaluation does not provide comparable conclusions as relying on human speech for the particular task of automatic speech recognition (ASR) and the FLEURS domain (see Section 4).
@@
We commission human recordings for the part of the Belebele dataset that is not covered by existing FLEURS recordings, as well as for elements of Belebele that do not exist in FLEURS (i.e. questions and answers). Recording participants must be native speakers of the languages they record. They must have an impeccable grasp of the conventions used in their respective languages for the narration of texts. The three tasks that participants are asked to perform are: (1) Read aloud and record the text passages provided (from FLORES-200); (2) Read aloud and record the provided written questions; (3) Read aloud and record the provided written answers.
For the task, we provide the participants with (a) the text of the sentences to be recorded in TSV format (the number of passages may differ from language to language), (b) the written questions (900 per language, and (c) the written answer options (3,600 per language). Additional details on the recording guidelines provided to annotators are reported in the appendix B.
We verify the quality of the recordings by randomly selecting 270 recordings (30% of sample size) and ensuring that the recordings do not contain background or ambient noise and that the voices of the participants are clearly audible.
@@
To obtain ASL sign recordings, we provide translators of ASL and native signers with the English text version of the sentences to be recorded. The interpreters are then asked to translate these sentences into ASL, create glosses for all sentences, and record their interpretations into ASL one sentence at a time. The glosses are subjected to an additional quality check by expert annotators to harmonize the glossing format.
To harmonize the recording conditions and eliminate visual bias, the videos are recorded against plain monochrome backgrounds (e.g., white or green), and signers are requested to wear monochrome upper body clothing (e.g., black). All videos are captured in 1920x1080p resolution with all of the signing space covered in FOV. The recordings are done in 60 frames per second to address most of the motion blur that happens during signing.
@@
The final dataset is composed of 75 languages (74 in speech, 1 in sign). Each of the languages’ respective subsets includes 2,000 utterances organized in 488 distinct passages, 900 questions, and 4 multiple choice answers per question. For our recorded data (the red portion of Figure 1 plus questions and answers), we have one audio file or two per sentence, depending on the number of available participants (one participant only in 23 languages, and two participants in 51 languages). When two speakers are available, we request that one should represent a higher-pitch range, and the other a lower-pitch range for each passage. Additional details are available in Table 4.
@@ 3 Experiments
We evaluate 2M-Belebele, and compare performance across modalities. Our comparison is limited in number of systems and combination of modalities. 2M-Belebele offers the opportunity to check multimodal comprehension by combining speech/text/sign passages; questions and answers. In our case, we only provide results for entire text passages, questions and answers and speech or sign passages, text questions and answers. The full set of experiments with multiple models is out of the scope of this paper, which aims at unblocking such experimentation by open-sourcing the dataset itself.
@@
| Belebele | Llama-3 | 70B | 128K | - | - | - | 94.8 | - |
| 2M-Belebele | Whisper + Llama-3 | 70B | 128K | - | - | - | 94.4 | - |
| 2M-Belebele | SeamlessM4T + Llama-3 | 70B | 32K | - | - | - | 93.5 | |
@@
We use the speech section of the 2M-Belebele dataset to evaluate the speech comprehension task with a cascaded system consisting of first speech recognition (ASR) using the Whisper-large-v3 model (Radford et al., 2022) (hereinafter, Whisper) and SeamlessM4T (corresponding to SeamlessM4T-Large v2) model Seamless-Communication et al. (2023) feeding into Llama-3 222https://ai.meta.com/blog/meta-llama-3/ LLM. We also provide results with a unified system Spiritlm (Nguyen et al., 2024), which is a multimodal language model that freely mixes text and speech. Since the size of this model is 7B and is based on Llama-2, we also add a comparison to the Llama-2 model. We compare these results with Llama-3 and Llama-3-chat using the Belebele text passage as input. For these systems, we report the results in English in 5-shot in-context learning. For the best performing system (Whisper + Llama-3) we report zero-shot results on 39 languages at the intersection of Whisper, Seamless and 2M-Belebele (see column ASR in Table 4 in the Appendix A).
@@
We use the same few-shot template as Bandarkar et al. (2023), the few-shot examples are taken randomly from the English training set and they are prompted as text format to the model.
Zero-shot Evaluation.
Again, we use the same evaluation strategy as in Bandarkar et al. (2023). Spiritlm is not available in chat mode.
For 5-shot and zero-shot settings, we report the averaged accuracy over 3 runs333Random seeds: 0, 1, 2..
@@ Results.
2M-Belebele accuracy results per language compared to Belebele are shown in Figure 2. Differences in speech and text vary slightly depending on the languages. Low-resource languages have a greater variation between text and speech Belebele. The ten languages with the largest gap are: Burmese, Maltese, Assamese, Telugu, Javanese, Tajik, Bengali, Shona, Eastern Panjabi, Yoruba, Gujarati.
Table 2 reports the summary of the results. The English drop from direct text to speech task does not vary much between 5-shot and zero-shot strategies, being slightly higher in the 5-shot setting. Table 2 additionally reports English results for Spiritlm, a direct multimodal model. One of the reasons Spiritlm may be performing worse is that 5-shot examples are in text, while the passage on the asked question is in speech.
@@
We know from previous large-scale translation attempts (Albanie et al., 2021; Müller et al., 2022) that models struggle to generalize over both individuals/appearance and large domain of discourse. Compared to speech and text models, sign language models suffer from having to learn generalized representations from high-dimensional inputs, i.e. videos, without overfitting to limited training dataset. Previous attempts have been made to create a more generalizable abstraction layer in the form of subunits (Camgoz et al., 2020), similar to phonemes for speech, which achieved promising results on a translation task with a small discourse domain. However, this work is yet to be applied to large discourse domain translation tasks. The best results in the FLORES-200 domain have been achieved with close models that are not available (Tanzer, 2024) and their performance is below 10 BLEU. This is why we do not even report results with open models (Rust et al., 2024) that do not have a chance to perform above chance in the final reading comprehension dataset.
We believe that the release of this new dataset with the additional gloss annotation will help in training models that generalize over individuals better and improve large-scale sign language translation performance.
4 Ablation study: Synthetic extension in speech evaluation datasets
For our experiments, we generate a synthetic copy of the FLEURS dataset using the MMS TTS (Pratap et al., 2024) system on the FLEURS transcripts. Then, we benchmark state-of-the-art models (Whisper, SeamlessM4T and MMS ASR) on both the original and synthetic datasets and analyze whether the conclusions remain consistent across both datasets. 444Note that we perform the study on the FLEURS languages that are covered by all MMS, Whisper and SeamlessM4T.
| Seamless | Whisper | MMS | |||||
| Hum | Syn | Hum | Syn | Hum | Syn | Inv | |
| Bengali | 14.1 | 21.1 | 114.7 | 105.8 | 14.6 | 25.0 | |
[…]
| Dutch | 9.9 | 20.0 | 8.5 | 19.7 | 12.4 | 28.3 | |
| English | 6.0 | 11.7 | 4.5 | 9.8 | 12.3 | 19.2 | |
[…]
| German | 8.5 | 13.9 | 5.2 | 12.3 | 10.5 | 20.8 | |
[…]
| Korean | 25.7 | 40.3 | 15.4 | 29.9 | 47.8 | 61.2 | |
[…]
| Russian | 10.2 | 18.6 | 5.6 | 17.4 | 18.8 | 34.3 | |
[…]
| Swedish | 15.4 | 20.1 | 11.3 | 19.1 | 21.0 | 27.8 | |
| Telugu | 27.4 | 28.0 | 132.2 | 133.9 | 24.2 | 27.8 | |
| Thai | 127.8 | 135.5 | 104.0 | 121.3 | 99.8 | 99.9 | |
| Turkish | 18.6 | 23.0 | 8.4 | 16.5 | 19.2 | 30.3 | |
| Ukrainian | 15.0 | 23.5 | 9.8 | 21.8 | 18.1 | 34.7 | |
| Vietnamese | 16.0 | 20.1 | 10.2 | 14.2 | 25.8 | 25.3 | |
@@ 5 Conclusions
The 2M-Belebele dataset555Scripts to build the 2M-Belebele dataset are available in the Belebele github https://github.com/facebookresearch/belebele allows to evaluate natural language comprehension in a large number of languages, including ASL. 2M-Belebele is purely human-made and covers the Belebele passages, questions, and answers for 75 languages: 74 in the speech modality and 1 in the sign modality.
@@
Our speech annotations do not have the entire set completed with two annotators. Due to the high volume of the dataset, not every recording has been thoroughly verified. Some of the languages in 2M-Belebele are low-resource languages, which pose a challenge in sourcing professionals to record. Therefore, some of the audios were recorded in home settings and may contain minor background noise, static noise, echoes, and, occasionally, the speech could be slightly muffled or soft. All annotators are native speakers of the target language, but they may have regional accents in their speech, and their personal speech styles may be present in the audio as well. However, the mentioned imperfections should not affect intelligibility; all the recordings can be clearly understood by human standards.
@@
Speakers and signers were paid a fair rate. Our recorded data reports self-identified gender by participant.
@@ Appendix B Annotation Guidelines
Number formats can vary from language to language; it is important to follow the pronunciation rules in your language. Here are some general guidelines and examples: Decimal numbers: Read the whole part of the number as a whole number and then individually read every number after the decimal point. For example, in English, the decimal number 3.14 should be read as "three point one four." Different languages may have different rules, and you should follow the rules that are appropriate for your language. Cardinal numbers represent quantities or amounts. Ordinal numbers represent positions or ranks in sequential order and should be read with the appropriate suffix. For example, in English, the ordinal number 1st is read "first" (not "onest") and 5th is read "fifth" (not "fiveth"). Different languages may have different rules, and you should follow the rule that is appropriate for your language.
@@
Roman numerals are a collection of seven symbols that each represent a value: I = 1, V = 5, X = 10, L = 50, C = 100, D = 500, and M = 1,000. The can be pronounced in slightly different ways depending on the context, but they are never pronounced as individual letters. For example, in English, VIII in Henry VIII is pronounced "Henry the eighth", while Superbowl LVIII is pronounced "Superbowl fifty-eight", but they are never pronounced "Henry v i i i" or "Superbowl l v i i i". Different languages may have different rules, and you should follow the rules that are appropriate for your language. Punctuation marks: As a general rule, punctuation marks should not be pronounced, except quotation marks.
@@
For example, in English, punctuation marks such as periods, commas, colons, semicolons, question marks, and exclamation points are typically not pronounced. For example, the sentence. As a result of this, a big scandal arose. will be pronounced "As a result of this a big scandal arose" - not "As a result of this comma a big scandal arose period". However, in formal-register English (in the news, for example), a difference is made between content created by the news team and content that should be attributed to someone else by explicitly pronouncing quotation marks. For example, the news transcript The fighter said: "I am here to try to win this." will be pronounced: "The fighter said, quote, I am here to try to win this. End of quote." In this case, different languages may have different rules, and you should follow the rules that are appropriate for your language. Signs and symbols. Signs and symbols need to be pronounced as they would be heard in a speech-only setting. Attention should be paid: (a) to potential number or gender agreement (for example, in English, "40%" should be read as "forty percent" — not "forty percents") (b) to potential differences between the place of the sign or symbol in writing and in speech (for example, in English, the "$" sign should be read as "dollar" and should be read after the number it precedes; i.e. "$22" should be read as "twenty-two dollars" — not "dollars twenty-two") (c) to the way the sign or symbol gets expanded in speech (for example, in English, "Platform 9 ¾" should be read "platform nine and three quarters" — not "platform nine three quarters"). Similarly, 50 km/h would be pronounced "fifty kilometers per hour" — not "fifty kilometers hour"). Different languages may have different rules, and you should follow the rules that are appropriate for your language.
@@
Even the same language may have at least 2 different ways to pronounce foreign expressions of proper nouns: (a) one way is to try to approach the way they would sound in the foreign language from which they come (for example, in English, Louis in Louis XIV is pronounced "lewee" as it would be in French); (b) the other way is to pronounce them according to the rules of the adopting language (for example, in English, Louis in the City of St Louis is pronounced as in the English proper noun "Lewis")
@@
Abbreviations should be expanded as much as possible. However, it is suggested to refrain from expanding them if their expansion results in unnatural speech. For example, in English, abbreviations such as Dr. or etc. are pronounced "doctor" and "et cetera", respectively (not "d r" nor "e t c"). However, abbreviations such as AM or PhD are pronounced as a sequence of letters without being expanded ("a m" and "p h d", respectively - not "ante meridiem" nor "philosophy doctorate"). Different languages may have different conventions, and you should follow the conventions that are appropriate for your language.

### v2

@@ Abstract
We introduce the first highly multilingual speech and American Sign Language (ASL) comprehension dataset by extending Belebele. Our dataset covers 74 spoken languages at the intersection of Belebele and FLEURS, and one sign language (ASL). We evaluate 2M-Belebele dataset for both 5-shot and zero-shot settings and across languages, the speech comprehension accuracy is 2-3% average lower compared to reading comprehension.
2M-Belebele: Highly Multilingual Speech and American Sign Language Comprehension Dataset
Marta R. Costa-jussà, Bokai Yu, Pierre Andrews, Belen Alastruey, Necati Cihan Camgoz Joe Chuang, Jean Maillard, Christophe Ropers, Arina Turkantenko, Carleigh Wood FAIR, Meta {costajussa,bokai, mortimer,alastruey,neccam, joechuang,jeanmm,chrisropers,arinatur,carleighwood}@meta.com
@@ 1 Introduction
The recent Belebele benchmark is the first corpus that addresses text reading comprehension for a large number of languages following a multi-way parallel approach (Bandarkar et al., 2023). The entire Belebele text statistics are summarized in Table 1. Currently, there are no highly multilingual evaluation datasets for natural language understanding that cover either both speech and text, and/or ASL.
The outstanding performance of some MT and text-to-speech (TTS) models has enabled a rise in the number of works using synthetically generated training data. Furthermore, some recent works propose to also use synthetic data for evaluation; e.g., Üstün et al. (2024); Seamless-Communication et al. (2023); Nguyen et al. (2024); Nachmani et al. (2023). This strategy allows researchers to extend datasets to low-resource languages and to other modalities, such as speech. However, we prove that using synthetic data for evaluation does not provide comparable conclusions as relying on human speech for the particular task of automatic speech recognition (ASR) and the FLEURS domain (see Appendix D).
@@
We commission human recordings for the part of the Belebele dataset that is not covered by existing FLEURS recordings, as well as for elements of Belebele that do not exist in FLEURS (i.e. questions and answers). Recording participants must be native speakers of the languages they record. They must have an impeccable grasp of the conventions used in their respective languages for the narration of texts. The three tasks that participants are asked to perform are: (1) Read aloud and record the text passages provided (from FLORES-200); (2) Read aloud and record the provided written questions; (3) Read aloud and record the provided written answers. For the task, we provide the participants with (a) the text of the sentences to be recorded in TSV format (the number of passages may differ from language to language), (b) the written questions (900 per language, and (c) the written answer options (3,600 per language). Additional details on the recording guidelines provided to annotators are reported in the appendix B. We verify the quality of the recordings by randomly selecting 270 recordings (30% of sample size) and ensuring that the recordings do not contain background or ambient noise and that the voices of the participants are clearly audible.
@@
To obtain ASL sign recordings, we provide translators of ASL and native signers with the English text version of the sentences to be recorded. The interpreters are then asked to translate these sentences into ASL, create glosses for all sentences, and record their interpretations into ASL one sentence at a time. The glosses are subjected to an additional quality check by expert annotators to harmonize the glossing format. To harmonize the recording conditions and eliminate visual bias, the videos are recorded against plain monochrome backgrounds (e.g., white or green), and signers are requested to wear monochrome upper body clothing (e.g., black). All videos are captured in 1920x1080p resolution with all of the signing space covered in FOV. The recordings are done in 60 frames per second to address most of the motion blur that happens during signing.
@@
The final dataset is composed of 75 languages (74 in speech, 1 in sign). Each of the languages’ respective subsets includes 2,000 utterances organized in 488 distinct passages, 900 questions, and 4 multiple choice answers per question. For our recorded data (the red portion of Figure 1 plus questions and answers), we have one audio file or two per sentence, depending on the number of available participants (one participant only in 23 languages, and two participants in 51 languages). When two speakers are available, we request that one should represent a higher-pitch range, and the other a lower-pitch range for each passage. More details are available in Appendix A.
@@ 3 Experiments
We evaluate 2M-Belebele, and compare performance across modalities. Our comparison is limited in number of systems and combination of modalities. 2M-Belebele offers the opportunity to check multimodal comprehension by combining speech/text/sign passages; questions and answers. In our case, we only provide results for entire text passages, questions and answers and speech passages, text questions and answers. A more comprehensive set of experiments is out of the scope of this paper, which aims at unblocking such experimentation by open-sourcing the dataset itself.
@@
| Belebele | Llama-3 | 70B | 128K | 84.9 | 97.4 | 94.9 | 94.8 | 84.7 |
| 2M-Belebele | Whisper + Llama-3 | 70B | 128K | 77.1 | 89.7 | 71.8 | 94.4 | 76.6 |
| 2M-Belebele | SeamlessM4T + Llama-3 | 70B | 128K | 81.7 | 94.9 | 92.7 | 93.5 | 81.4 |
@@
We use the speech section of the 2M-Belebele dataset to evaluate the speech comprehension task with a cascaded system consisting of first speech recognition (ASR) using the Whisper-large-v3 model (Radford et al., 2022) (hereinafter, Whisper) and SeamlessM4T (corresponding to SeamlessM4T-Large v2) model Seamless-Communication et al. (2023) feeding into Llama-3 222https://ai.meta.com/blog/meta-llama-3/. We also provide results with a unified system Spiritlm (Nguyen et al., 2024), which is a multimodal language model that freely mixes text and speech. Since the size of this model is 7B and is based on Llama-2, we also add a comparison to the Llama-2 model. We compare these results with Llama-3 and Llama-3-chat using the Belebele text passage as input. For these systems, we report the results in 5-shot in-context learning and zero-shot on 39 languages at the intersection of Whisper, SeamlessM4T and 2M-Belebele (see Appendix A).
Zero-shot Evaluation.
We use the same evaluation strategy as in Bandarkar et al. (2023). Spiritlm is not available in chat mode.
@@
The few-shot examples are taken randomly from the English training set and they are prompted as text format to the model. Different from Bandarkar et al. (2023), we do not pick the answer with the highest probability but directly assess the predicted letter of the answer.
For 5-shot and zero-shot settings, our instruction prompt is as follows “Given the following passage, query, and answer choices, output the letter corresponding to the correct answer. Do not write any explanation. Only output the letter within A, B, C, or D that corresponds to the correct answer.” and we report the averaged accuracy over 3 runs333Random seeds: 0, 1, 2..
@@ Results.
2M-Belebele accuracy results per language compared to Belebele are shown in Figure 2. Differences in speech and text vary slightly depending on the languages. Low-resource languages have a greater variation between text and speech Belebele. The ten languages with the largest gap are: Burmese, Maltese, Assamese, Telugu, Javanese, Tajik, Bengali, Shona, Eastern Panjabi, Yoruba, Gujarati. Table 2 reports the summary of the results. The English drop from direct text to speech task does not vary much between 5-shot and zero-shot strategies, being slightly higher in the zero-shot setting (coherently with previous Llama-3 results that show better performance in zero-shot in other tasks444https://ai.meta.com/blog/meta-llama-3-1/ and https://ai.meta.com/blog/meta-llama-3/). When comparing speech and text comprehension, we observe that speech decreases performance in about 2-3% average across languages depending on the setting. Table 2 reports English results for Spiritlm, a direct multimodal model. One of the reasons Spiritlm may be performing worse is that 5-shot examples are in text, while the passage on the asked question is in speech.
@@
We know from previous large-scale translation attempts (Albanie et al., 2021; Müller et al., 2022) that models struggle to generalize over both individuals/appearance and large domain of discourse. Compared to speech and text models, sign language models suffer from having to learn generalized representations from high-dimensional inputs, i.e. videos, without overfitting to limited training dataset. Previous attempts have been made to create a more generalizable abstraction layer in the form of subunits (Camgoz et al., 2020), similar to phonemes for speech, which achieved promising results on a translation task with a small discourse domain. However, this work is yet to be applied to large discourse domain translation tasks. The best results in the FLORES domain have been achieved with close models that are not available (Zhang et al., 2024). Trying Rust et al. (2024) as an open model did not perform above chance in the final reading comprehension dataset. However, we believe that the release of this new dataset with the additional gloss annotation will help in training models that generalize over individuals better and improve large-scale sign language translation.
@@ 4 Conclusions
The 2M-Belebele dataset5552M-Belebele dataset is freely available in Github https://github.com/facebookresearch/belebele and in HuggingFace https://huggingface.co/datasets/facebook/2M-Belebele allows to evaluate natural language comprehension in a large number of languages, including ASL. 2M-Belebele is purely human-made and covers the Belebele passages, questions, and answers for 75 languages: 74 in the speech modality and 1 in the sign modality.
@@
Our speech annotations do not have the entire set completed with two annotators. Due to the high volume of the dataset, not every recording has been thoroughly verified. Some of the languages in 2M-Belebele are low-resource languages, which pose a challenge in sourcing professionals to record. Therefore, some of the audios were recorded in home settings and may contain minor background noise, static noise, echoes, and, occasionally, the speech could be slightly muffled or soft. All annotators are native speakers of the target language, but they may have regional accents in their speech, and their personal speech styles may be present in the audio as well. However, the mentioned imperfections should not affect intelligibility; all the recordings can be clearly understood by human standards. Note that we are planning to release more languages as reported in Appendix C.
@@
In terms of compute budget, we estimate it as 47K Nvidia A100 hours by taking into account the product of following factors: number of languages (39), number of random seeds (3), number of GPUs required by model (8), number of experiment setups (5) and estimated number of hours per experiment (10).
Speakers and signers were paid a fair rate. Our recorded data reports self-identified gender by participant. Each of the speakers and signers signed a consent form agreeing on the dataset and its usage that they were participating in.
@@ Appendix B Annotation Guidelines
Number formats can vary from language to language; it is important to follow the pronunciation rules in your language. Here are some general guidelines and examples: Decimal numbers: Read the whole part of the number as a whole number and then individually read every number after the decimal point. For example, in English, the decimal number 3.14 should be read as ”three point one four.” Different languages may have different rules, and you should follow the rules that are appropriate for your language. Cardinal numbers represent quantities or amounts. Ordinal numbers represent positions or ranks in sequential order and should be read with the appropriate suffix. For example, in English, the ordinal number 1st is read ”first” (not ”onest”) and 5th is read ”fifth” (not ”fiveth”). Different languages may have different rules, and you should follow the rule that is appropriate for your language.
@@
Roman numerals are a collection of seven symbols that each represent a value: I = 1, V = 5, X = 10, L = 50, C = 100, D = 500, and M = 1,000. The can be pronounced in slightly different ways depending on the context, but they are never pronounced as individual letters. For example, in English, VIII in Henry VIII is pronounced ”Henry the eighth”, while Superbowl LVIII is pronounced ”Superbowl fifty-eight”, but they are never pronounced ”Henry v i i i” or ”Superbowl l v i i i”. Different languages may have different rules, and you should follow the rules that are appropriate for your language. Punctuation marks: As a general rule, punctuation marks should not be pronounced, except quotation marks.
@@
For example, in English, punctuation marks such as periods, commas, colons, semicolons, question marks, and exclamation points are typically not pronounced. For example, the sentence. As a result of this, a big scandal arose. will be pronounced ”As a result of this a big scandal arose” - not ”As a result of this comma a big scandal arose period”. However, in formal-register English (in the news, for example), a difference is made between content created by the news team and content that should be attributed to someone else by explicitly pronouncing quotation marks. For example, the news transcript The fighter said: ”I am here to try to win this.” will be pronounced: ”The fighter said, quote, I am here to try to win this. End of quote.” In this case, different languages may have different rules, and you should follow the rules that are appropriate for your language. Signs and symbols. Signs and symbols need to be pronounced as they would be heard in a speech-only setting. Attention should be paid: (a) to potential number or gender agreement (for example, in English, ”40%” should be read as ”forty percent” — not ”forty percents”) (b) to potential differences between the place of the sign or symbol in writing and in speech (for example, in English, the ”$” sign should be read as ”dollar” and should be read after the number it precedes; i.e. ”$22” should be read as ”twenty-two dollars” — not ”dollars twenty-two”) (c) to the way the sign or symbol gets expanded in speech (for example, in English, ”Platform 9 ¾” should be read ”platform nine and three quarters” — not ”platform nine three quarters”). Similarly, 50 km/h would be pronounced ”fifty kilometers per hour” — not ”fifty kilometers hour”). Different languages may have different rules, and you should follow the rules that are appropriate for your language.
@@
Even the same language may have at least 2 different ways to pronounce foreign expressions of proper nouns: (a) one way is to try to approach the way they would sound in the foreign language from which they come (for example, in English, Louis in Louis XIV is pronounced ”lewee” as it would be in French); (b) the other way is to pronounce them according to the rules of the adopting language (for example, in English, Louis in the City of St Louis is pronounced as in the English proper noun ”Lewis”)
@@
Abbreviations should be expanded as much as possible. However, it is suggested to refrain from expanding them if their expansion results in unnatural speech. For example, in English, abbreviations such as Dr. or etc. are pronounced ”doctor” and ”et cetera”, respectively (not ”d r” nor ”e t c”). However, abbreviations such as AM or PhD are pronounced as a sequence of letters without being expanded (”a m” and ”p h d”, respectively - not ”ante meridiem” nor ”philosophy doctorate”). Different languages may have different conventions, and you should follow the conventions that are appropriate for your language.